# sources.bib

### v1

@@ 4. Localizations of twisted Calabi–Yau algebras
We find ourselves now equipped to discuss examples arising from the ideas considered at the beginning of this section.
Example 4 (Localizations of polynomial rings over a field).
Recall that in this commutative case, any multiplicative subset of our algebra will provide a left and right denominator subset. As an immediate consequence of Theorem 4.7, we have that $k[t,t^{-1}]$ is Calabi–Yau of dimension 1. Further, since $k[t,t^{-1}]$ is $\mathbb{Z}$-graded with $\deg(t)=1$ and $\deg(t^{-1})=-1$, Theorem 3.2 immediately gives that $k[t,t^{-1}]$ is $\mathbb{Z}$-graded Calabi–Yau of dimension 1. Perhaps even more surprising is that if we consider the field of fractions of $k[t]$, Theorem 4.7 tells us that $k(t)$ is Calabi–Yau of dimension 1. Observe that $k(t)$ no longer carries an induced grading, since the multiplicative set we localize at contains non-homogeneous elements (alternatively, we can recall that a field can only have trivial $\mathbb{Z}$-grading). Of course, one can construct many similar examples by considering different localizations of the $n$-dimensional Calabi–Yau algebras $k[t_{1},\dots,t_{n}]$.
@@
In leaving the commutative setting, the question of finding suitable sets at which to localize becomes more difficult. Here is one potentital strategy. Recall that an element $x$ of a ring $R$ is said to be normal if $xR=Rx$. In particular, the set of normal elements of a ring will generate a multiplicatively closed right and left permutable subset. If $S$ is the set of normal elements of a ring $R$, then so long as $S$ is left and right reversible, $S$ will provide a left and right denominator subset of $R$. In the case where $R$ is a domain, this comes for free, and we find applications in the following example.
Example 5 (Localizations of Artin–Schelter regular algebras).
Closely related to the study of Calabi–Yau algebras is that of Artin–Schelter regular algebras, a definition of which can be found in [rogalski2023artinschelter], for example. Capturing this relationship is Lemma 1.2 of [Reyes_2014], which asserts that connected $\mathbb{N}$-graded Calabi–Yau algebras are Artin–Schelter regular. Thus, in light of Theorem 4.7 and the previous example, we can expect to find a large number of $\mathbb{Z}$-graded Calabi–Yau algebras arising as localizations of Artin–Schelter regular algebras, provided we have a suitable left and right denominator subset.
Recall that all known examples of Artin–Schelter regular algebras are domains, and that it is conjectured that this is always the case. In light of the above discussion, localizing any (known) Artin–Schelter regular algebra at its set of homogeneous normal elements will provide an example of a $\mathbb{Z}$-graded Calabi–Yau algebra. Another choice of a left and right denominator set is given by $\{1,f,f^{2},\dots\}$ for $f$ a homogeneous normal element. Further, given that the algebra has finite GK-dimension, we could also localize at the set of all nonzero elements, thus giving rise to more examples of Calabi–Yau division algebras.

### v2

@@ 4. Localizations of twisted Calabi–Yau algebras
We are now equipped to discuss examples arising from the ideas considered at the beginning of this section.
Example 4.8 (Localizations of polynomial rings over a field).
Let $A=k[t_{1},\dots,t_{n}]$. Because $A$ is commutative, any multiplicative subset will provide a left and right denominator subset. As a consequence of Theorem 4.7, we have that $k[t,t^{-1}]$ is Calabi–Yau of dimension 1. Further, since $k[t,t^{-1}]$ is $\mathbb{Z}$-graded with $\deg(t)=1$ and $\deg(t^{-1})=-1$, Theorem 3.2 gives that $k[t,t^{-1}]$ is $\mathbb{Z}$-graded Calabi–Yau of dimension 1. Furthermore, by Theorem 4.7 the field of fractions $k(t)$ is Calabi–Yau of dimension 1. Observe that $k(t)$ no longer carries an induced grading, since the multiplicative set we localize at contains non-homogeneous elements. Alternatively, we can recall that a field can only have trivial $\mathbb{Z}$-grading. One can construct many similar examples by considering different localizations of the $n$-dimensional Calabi–Yau algebras $k[t_{1},\dots,t_{n}]$.
@@
Outside the commutative setting, the question of finding suitable sets at which to localize becomes more difficult. Here is one potential strategy. Recall that an element $x$ of a ring $R$ is said to be normal if $xR=Rx$. In particular, the set of normal elements of a ring will generate a multiplicatively closed right and left permutable subset. If $S$ is the set of normal elements of a ring $R$, then so long as $S$ is left and right reversible, $S$ will be a left and right denominator subset of $R$. In the case where $R$ is a domain, this comes for free, and we find applications in the following example.
Example 4.9 (Localizations of Artin–Schelter regular algebras).
By Theorem 4.7, we can expect to find a large number of $\mathbb{Z}$-graded Calabi–Yau algebras arising as localizations of Artin–Schelter regular algebras, provided we have a suitable left and right denominator subset.
All known examples of Artin–Schelter regular algebras are domains, and it is conjectured that this is always the case. In light of the above discussion, localizing any (known) Artin–Schelter regular algebra at its set of homogeneous normal elements is an example of a $\mathbb{Z}$-graded Calabi–Yau algebra. Another choice of a left and right denominator set is given by $\{1,f,f^{2},\dots\}$ for $f$ a homogeneous normal element. Further, given that the algebra has finite GK-dimension, we could also localize at the set of all nonzero elements, thus giving rise to more examples of Calabi–Yau division algebras.